The lengths of the sides of a triangle are consecutive integers. Half of the perimeter is 14 more than the length of the longest side. Find the perimeter.
Answer by John10(297)

The lengths of the sides of a triangle are consecutive integers. Half of the perimeter is 14 more than the length of the longest side. Find the perimeter.
---------------------------
Let's x, x+1, and x + 2 are the "consecutive integer" sides of the triangle
Half of the perimeter is 14 more than the length of the longest side:
P/2 = 14 + (x + 2)
P = 28 + 2(x + 2)
P = 28 + 2x + 4
P = 2x + 32
Have a great one! John10